# SciClaims: An End-to-End Generative System for Biomedical Claim Analysis

### v1

@@ 5.2.1 Claim quality evaluation
In this phase, we evaluate the quality of the claims generated by each method. We devise a questionnaire with eight yes/no questions (see Table 3) to capture the desired properties in a correct claim and ask the judge model to answer it. The first question requires context from the source paragraph, while the remaining questions focus solely on the claim. Correct claims need to receive a Yes to all questions. Table 4 shows that our LLM-based system outperforms all other methods, scoring 15 points higher than the second-best, the sentence tokenizer.
Id Question
Q1 Is the claim grounded by the original text?
Q2 Is the claim grammatically correct?
Q4 Is the claim precise and specific rather than vague?
Q6 Is the claim concise without losing essential information?
Q8 Would verifying the claim add value to public knowledge?
Table 3: Questions asked to the judge to evaluate the quality of the generated claims.
@@
Generation Module Verification Module Label Accuracy Not NEI (%) Label Accuracy Not NEI (%) System Score Time/doc (secs.)
@@
Pan et al MultiVerS 0.7591 2.31 0.5714 11.90 0.0180 2.17
Wright et al MultiVerS 0.5015 5.53 0.4268 8.35 0.0482 14.25
@@
Pan et al SciClaims 0.6568 39.60 0.619 71.43 0.0195 7.61
Wright et al SciClaims 0.6397 69.44 0.7098 74.66 0.0801 65.28
@@
Table 5: Verification and full-system evaluation (phase 3) results with a judge model (Qwen2.5 72B AWQ). The processing time per document is the average time in seconds that each system needs to process a single document.
@@ 5.2.2 Document retrieval evaluation
Claims All Claims Correct Claims R@1 R@3 R@5 R@1 R@3 R@5 Sentence tokenizer 0.5591 0.7205 0.7795 0.6265 0.7390 0.7871 Pan et al 0.2970 0.4653 0.5248 0.6429 0.7143 0.7143 Wright et al 0.5860 0.7474 0.8038 0.7113 0.8365 0.8726 Noun Phrases Gen 0.4634 0.6456 0.7013 0.6286 0.7619 0.8095 SciClaims 0.5045 0.6645 0.7135 0.5257 0.6938 0.7344 SciClaims+CDP 0.5146 0.6940 0.7466 0.5944 0.7448 0.7867 SciClaims+CDP+CR 0.4727 0.6250 0.6777 0.5102 0.6675 0.7107
@@ 5.2.3 Claim verification and full-system
SciClaims demonstrates superior claim verification accuracy, outperforming the fine-tuned MultiVerS Wadden et al. (2022) model despite operating in a zero-shot setting. As shown in Table 5, SciClaims consistently produces more accurate labels across all claim generation methods, except Pan et al. (2021). Notably, SciClaims returns significantly fewer NEI labels than MultiVerS, providing greater value to users by delivering more definitive SUPPORT or REFUTE labels. The unusually high accuracy of Pan et al. likely stems from its excessive NEI labels, which simplify label selection. Furthermore, the SciClaims+CDP+CR system achieves the highest overall performance, correctly labeling 50% of generated claims, outperforming the next-best MultiVerS-based system by over five points.
@@ 6 Conclusion
Built on state-of-the-art LLMs, SciClaims streamlines claim extraction, evidence retrieval, and verification, outperforming previous methods in both claim extraction and verification without additional fine-tuning. It also delivers user-friendly explanations in natural language, ensures competitive throughput in terms of processing time per input document, and leverages its underlying LLM’s context window to handle larger documents than previous encoder-based approaches. Moreover, SciClaims is also frugal hardware-wise. Major pharmaceutical companies like Sanofi and scientific literature review platforms such as DistillerSR are currently exploring SciClaims for real-world applications including patent infringement detection.
@@ A.9 Phase 3 evaluation with judge model
Appendix B Screenshots
Figure 2: Screenshot of the SciClaims Demo examples panel.
Figure 3: Screenshot of the SciClaims Demo results.
Appendix C Additional metrics
Method Rouge1 Rouge2 RougeL BertScore Similarity Score
Sentence tokenizer 0.3313 0.198 0.3030 0.8001 0.7211
Pan et al 0.278 0.1334 0.2555 0.8152 0.6561
Wright et al 0.2525 0.0762 0.222 0.862 0.6475
Noun Phrases Gen 0.2567 0.0953 0.2293 0.8753 0.6693
SciClaims 0.3387 0.1896 0.3084 0.8204 0.725
Table 7: Results in Claim Generation in SciFact
Method Rouge1 Rouge2 RougeL BertScore Similarity Score
Sentence tokenizer 0.2846 0.1332 0.268 0.8173 0.6549
Pan et al 0.3722 0.1801 0.3477 0.8816 0.7147
Wright et al 0.361 0.1559 0.3359 0.8761 0.7023
Noun Phrases Gen 0.3383 0.141 0.3146 0.8844 0.7018
SciClaims 0.3448 0.1722 0.3219 0.8244 0.7028
Table 8: Results in Claim Generation in FEVER
Method Rouge1 Rouge2 RougeL BertScore Similarity Score
Sentence tokenizer 0.2905 0.1333 0.2744 0.8587 0.6667
Pan et al 0.2796 0.134 0.2569 0.8155 0.7128
Wright et al 0.3375 0.1384 0.3188 0.8803 0.6786
Noun Phrases Gen 0.3213 0.1326 0.302 0.8881 0.6889
SciClaims 0.3533 0.1708 0.3345 0.8802 0.7184
Table 9: Results in Claim Generation in RoSE

### v2

@@ 5.2.1 Claim quality evaluation
In this phase, we evaluate the quality of the claims generated by SciClaims and compare it to other methods, using the same baselines mentioned in Section 5.1.
@@
Claims All Claims Correct Claims R@1 R@3 R@5 R@1 R@3 R@5 Sentence tokenizer 0.5591 0.7205 0.7795 0.6265 0.7390 0.7871 Pan et al. (2021) 0.2970 0.4653 0.5248 0.6429 0.7143 0.7143 Wright et al. (2022) 0.5860 0.7474 0.8038 0.7113 0.8365 0.8726 Noun Phrases Gen 0.4634 0.6456 0.7013 0.6286 0.7619 0.8095 SciClaims 0.5045 0.6645 0.7135 0.5257 0.6938 0.7344 SciClaims+CDP 0.5146 0.6940 0.7466 0.5944 0.7448 0.7867 SciClaims+CDP+CR 0.4727 0.6250 0.6777 0.5102 0.6675 0.7107
We devised a questionnaire consisting of eight yes/no questions (see Appendix B) to capture the desired properties in a correct claim and ask the judge model to answer it. The first question requires context from the source paragraph, while the remaining questions focus solely on the claim. Correct claims need to receive a Yes to all questions. Table 3 shows that our LLM-based system outperforms all other methods, scoring 15 points higher than the second-best, the sentence tokenizer.
@@
Extraction Module Verification Module Label Accuracy Not NEI (%) Label Accuracy Not NEI (%) System Score Time/doc (secs.)
@@
Pan et al. (2021) MultiVerS 0.7591 2.31 0.5714 11.90 0.0180 2.17
Wright et al. (2022) MultiVerS 0.5015 5.53 0.4268 8.35 0.0482 14.25
@@
Pan et al. (2021) SciClaims 0.6568 39.60 0.619 71.43 0.0195 7.61
Wright et al. (2022) SciClaims 0.6397 69.44 0.7098 74.66 0.0801 65.28
@@
Table 5: Verification evaluation (phase 3) results with a judge model (Qwen2.5 72B AWQ). The table shows the label accuracy of each combination of extraction and verification module. The Correct Claims column counts only those considered correct in the Phase 1 questionnaire. Not NEI represents the proportion of claim–evidence pairs labeled as SUPPORT or REFUTE, rather than NEI (Not Enough Information). The last two columns summarize system-level performance: Processing Time per Document is the average time (in seconds) each system takes to process a single document, and System Score reflects the ratio of correctly labeled and well-formed claims across the system.
@@ 5.2.3 Claim verification and full-system
The verification task involves predicting the veracity of claim-evidence pairs, classifying each as SUPPORT, REFUTE, or NOT ENOUGH INFORMATION (NEI). Thus we evaluate our system’s label accuracy by asking to the judge model whether the assigned label is correct. We compare SciClaims results with MultiVerS Wadden et al. (2022).
SciClaims demonstrates superior claim verification accuracy, outperforming the fine-tuned MultiVerS Wadden et al. (2022) model despite operating in a zero-shot setting. As shown in Table 5, SciClaims consistently produces more accurate labels across all claim generation methods, except Pan et al. (2021). Notably, SciClaims returns significantly fewer NEI labels than MultiVerS, providing greater value to users by delivering more definitive SUPPORT or REFUTE labels. The unusually high accuracy of Pan et al. (2021). likely stems from its excessive NEI labels, which simplify label selection. Furthermore, the SciClaims+CDP+CR system achieves the highest overall performance, correctly labeling 50% of generated claims, outperforming the next-best MultiVerS-based system by over five points.
Table 5 also reports the average processing time per document for each system configuration. Systems using MultiVerS as the verification module tend to be the fastest overall, with the exception of those paired with claim extraction modules like Wright et al. (2022) and Noun Phrases Generation, which produce a high number of claims and thus increase runtime. Among all configurations, those using SciClaims+CDP+CR for claim generation—paired with either SciClaims or MultiVerS for verification—offer the best trade-off between predictive accuracy and time efficiency. Importantly, only the configuration that uses SciClaims as both the generation and verification module can run on a single 24GB GPU, making it uniquely suitable for real-world deployment.
5.3 Human evaluation
To complement our automated evaluation with LLMs, we conducted a human evaluation to assess the quality, relevance, and usability of SciClaims outputs. Five NLP-experts independently reviewed a sample of 154 PubMed documents processed by the system. The annotation tasks were divided into three phases, following the same structure as our LLM-based evaluation (see Section 5.2).
Claim quality evaluation. Annotators were shown an input paragraph and one randomly selected claim extracted by SciClaims. They answered the same eight binary questions used in the LLM-based evaluation (see Appendix B), assessing conciseness, precision, and other key dimensions. Table 6 shows that human judgments closely align with the judge model, with 70.5% of the claims meeting all eight criteria.
Claim Quality Score Rationale Quality Score Q1 (grounding) 0.9048 Q2 (grammar) 0.9810 Q3 (completeness) 0.9810 RQ1 (justification) 0.8571 Q4 (precision) 0.8667 RQ2 (relevance) 0.9206 Q5 (relevance) 0.8952 RQ3 (completeness) 0.7937 Q6 (conciseness) 0.8571 Q7 (self-contained) 0.9238 Q8 (contribution) 0.8857 Correct Claims (%) 70.47 Correct Rationales (%) 74.61
Table 6: Claim and rationale evaluation of SciClaims by human annotators. Score indicates the overall ratio of ’yes’ responses to the question along the annotators.
Document retrieval evaluation. Annotators were presented with a claim and its corresponding retrieved paragraph from SciClaims. They were asked whether the retrieved evidence provided sufficient information to verify the claim. In 60% of cases, annotators judged the paragraph as informative enough for claim verification, indicating moderate effectiveness of the retrieval module in real-world scenarios.
Claim verification evaluation. In this phase, annotators were shown a claim, its evidence, the system’s predicted label, and the corresponding rationale. They rated the label’s accuracy on a five-point scale, from 1 (completely inaccurate) to 5 (highly accurate). SciClaims achieved a strong average score of 4.40, reflecting high alignment between system predictions and human judgment. Additionally, annotators evaluate the quality of the generated rationale using three yes/no questions: (RQ1) Is the label justified by the rationale? (RQ2) Does the rationale focus on relevant information? (RQ3) Does the rationale provide enough context to understand the label?. As shown in Table 6, the responses from annotators indicate high satisfaction with the relevance and justification of rationales, though they also noted that completeness could be improved.
@@ 6 Conclusion
We presented SciClaims, a practical, end-to-end system for scientific claim analysis in the biomedical domain. Through an interactive web-based interface, users can extract, verify, and explore scientific claims with evidence-backed rationales, all powered by LLMs and a curated biomedical corpus. SciClaims is designed for usability and speed, and it is already being explored in real-world settings such as pharmaceutical patent analysis and systematic literature reviews. By combining explainable outputs with efficient infrastructure, it provides a robust tool for researchers, clinicians, and analysts who need to validate scientific information quickly and reliably. The system is openly accessible and ready for demonstration, offering an engaging experience to showcase the capabilities of modern LLM-based scientific reasoning.
Acknowledgments
The authors gratefully acknowledge the EU Large Language Models for EU (LLMs4EU) project (DIGITAL-20234-AI-06-LANGUAGE-01) and the HORIZON FAIR to Adapt to Climate Change (FAIR2Adapt) grant (agreement 101188256) for their support during the development of this work.
@@ A.9 Phase 3 evaluation with judge model
Appendix B Claim Quality Questionnaire
Id Question
Q1 Is the claim grounded by the original text?
Q2 Is the claim grammatically correct?
Q4 Is the claim precise and specific rather than vague?
Q6 Is the claim concise without losing essential information?
Q8 Would verifying the claim add value to public knowledge?
Table 7: Questions asked to the LLM judge and human annotators to evaluate the quality of the generated claims.